OpenStudy (anonymous):

solve the proportion. x/5=16/10

OpenStudy (anonymous):

x=16/10 *5 x=16/2=8 x=8

jhonyy9 (jhonyy9):

x/5 =16/10 10x=80 x=80/10 =8 x=8

jhonyy9 (jhonyy9):

but if you see 16/10 you can simplify with 2 and hence you get 8/5 in this way resulted that x/5 = 8/5 --- from this resulted that x=8